Docstoc

ISCC November meeting 2009

Document Sample
ISCC November meeting 2009 Powered By Docstoc
					  ISOLDE scientific coordinator’s report
             ISCC meeting, November 16, 2009



A. Herlert, CERN PH-SME-IS


             Running period 2009
             Status of experiments
             New ISOLDE web page
                 ISOLDE experiments and requests 2009


             INTC February 2009
                  • 10 new experiments: IS484 - IS493
                  • 363 approved RIB shifts
                  • 83 experiments after February meeting
             Beam requests 2009
                  • 11 completed experiments
                  • 51 experiments out of 72 requested beam (627 shifts)
                  • Requests for REX-ISOLDE 273 shifts
             INTC May 2009
                  • 1 new experiment: IS494
                  • 21 approved RIB shifts



A. Herlert                                                     ISCC November 16, 2009
                ISOLDE schedule 2009 (Apr 20 - Nov
                              23)
   Requested shifts:
    • 662 in total
       • 96 different beams
       • 627 for 51 experiments

   Scheduled shifts (~70%):
     • 463 in total
       • 62 different beams
       • 437 for 39 experiments

   REX-ISOLDE:
    • 273 shifts requested
       • 140 shifts scheduled
       • 9 experiments (8 runs)
       • additional WITCH runs
         (need REXTRAP)
                                  GPS   HRS

A. Herlert                                     ISCC November 16, 2009
                             Main problems in 2009 (I)
 Wk 15: Startup problems -
 beam to users (COLLAPS) very
 much delayed (off-line tests possible)
Wk 17/18: Vacuum problems, target
performance not ok, power cut,
controls problems ... debugging run
 Wk 19: Target change delayed
 due to PISOMOVE problems,
 11Li yields not sufficient but

 detector tests possible (IS417)
Wk 20: PSB injection septum broken
IS457 has to run with limited intensity
 Wk 22/23: REXLINAC 9-gap failure
 no beam for IS485
 Wk 25/26: YO target not ready for
 IS478 (cancelled). Rescheduled
 IS485 takes beam until several
 power cuts on June 18 eliminate
 ISOLDE operation (vacuum etc.)
                                          GPS   HRS

A. Herlert                                               ISCC November 16, 2009
                            Main problems in 2009 (II)

Wk 31: IS443 stopped due to HV
tripping ... some data obtained

 Wk 35: LINAC2 vacuum probem,
 IS480 stopped shortly after start of
 run - no data


 Wk 42: REXEBIS cathode broken,
 IS430 resumes measurements after
 very fast repair

 Wk 44: HV problems, wrongly set
 extraction electrode and ripple on
 HV (missing ground) - no beam
 for IS413

 Wk 47: HRS front-end blocked, no
 target change possible - IS433
 cannot take beam from HRS but
 from GPS as scheduled
                                        GPS    HRS

A. Herlert                                               ISCC November 16, 2009
                Accelerator schedule 2010 (draft)

 ISOLDE dates:
 protons from PSB April 12
 physics start ISOLDE April 29
 protons stop November 22
 (i.e. 29.5 weeks for physics
 as compared with 31 weeks
 in 2009)

 Longer shutdown period
 due to new HRS front-end
 and new vacuum system

 Several long periods of MD
 partly with parasitic beam to
 ISOLDE (not guaranteed)

A. Herlert                                      ISCC November 16, 2009
              Follow up: Reallocation of floor space
                      for experiments 2008


                                    ISOLTRAP
                                               MISTRAL




                                                                     MINIBALL
        ECR



                            WITCH                 ?



A. Herlert                                         ISCC November 16, 2009
              Follow up: Reallocation of floor space
                    for experiments - present


                                    ISOLTRAP
                                               MISTRAL




                                                                     MINIBALL
         EC



                            WITCH
                                                                                b




A. Herlert                                         ISCC November 16, 2009
                     Status of experiments - ECR


              Aug 2009
             ECR removal and new GHM experiment




                          ECR experiment before removal
A. Herlert                                                ISCC November 16, 2009
             Status of experiments - ECR removed

                Aug 2009                   Parts to be shipped
                                           to Grenoble and
                                           GANIL

                                           Some equipment
                                           belongs to ISOLDE
                                           and can be used
                                           by other users and
                                           experiments

                                           GHM collection
                                           chamber removed,
                                           new chamber to
                                           be designed and
                                           installed
A. Herlert                                   ISCC November 16, 2009
                        Status of experiments - EC

             Aug 2009




                    New experiment: On-line emission channeling
A. Herlert                                                  ISCC November 16, 2009
               Status of experiments - MISTRAL


• Clean-up has started

• RFQ buncher and cooler of
  MISTRAL removed (shipped
  to University Mainz)
• Looking into the removal of
  22t magnet
  difficult!




                                Oct 2009
A. Herlert                                   ISCC November 16, 2009
                Status of experiments - Beta-NMR

Two PhD students:
      Hans Tornqvist (CERN - supervisor Thomas Nilsson)
      Christophe Sotty (Orsay - supervisor Georgi Georgiev)

setup progressing well       Nov 2009

mailing list and
web page available

Initial safety discussion
to be done soon

Support from W.-D. Zeitz

Tilted foil polarization
First test in week 48 with
stable Ne beam from EBIS
A. Herlert                                                ISCC November 16, 2009
                   Status of experiments - CRIS


• Installation ongoing

• Prototype experiment
  for new safety structure
• Ready for off-line beam
  during Summer
• First on-line run planned
  for October 2010




                              Nov 2009


A. Herlert                                        ISCC November 16, 2009
                  Status of experiments - WITCH


  • New shielding installed (to prevent
    influence of WITCH magenetic
    field on HCI beam steering)
    Not enough (scaling not working)
  • Vacuum system of WITCH and
    almost all electrodes of the
    spectrometer improved
  • Commissioning run in week 47
  • Permanent crew (post-doc
    Martin Breitenfeldt and technician
    Julien Thiboud)


                                          Nov 2009
A. Herlert                                           ISCC November 16, 2009
              Status of experiments - HV Platform

                                                       Nov 2009
• Can only be used for collections

• Acceleration stage not functional
• Need new transformer?

• Check vacuum system



      Renovation of
      HV platform needed




A. Herlert                                    ISCC November 16, 2009
                   Status of experiments - GLM

                                        Nov 2009
• Heavily used by solid state physics
  community

• New setup available and to be
  installed soon

• Improvement of concrete shielding
  planned




A. Herlert                                         ISCC November 16, 2009
                      Status of experiments - ASPIC


   Future of ASPIC under discussion
   • Former HMI Berlin (now HZ for material
     science) would like to pass on the
     equipment to the ISOLDE Collaboration
     until new user takes over responsibility
     (no additional funds expected)
   • Kay Potzger showed interest, tries
     to obtain German BMBF funds to revive research program:
     quote from a recent email: "[...] I decided to write the application for re-installation
     of ASPIC. My boss (J. Fassbender) in FZD finally agrees and supports the project
     by involving our workshop, i.e. the new ASPIC chamber will be set up at FZD.
     Consequently, we will be main users but share the system with all the other
     groups from ISOLDE (unlike old ASPIC). [...]"


A. Herlert                                                                 ISCC November 16, 2009
                      New ISOLDE web page

 ALL Files and content transferred
 by Albert Deh Hwa Taureg
 (August 2009, "enfants de
 fonctionnairs" program at CERN)

 IT Security issues checked

 New web page located at:
 http://isolde.web.cern.ch/ISOLDE/
 isoldenewweb

 Many files need to be updated
 and new pages have to be
 created (e.g. for safety)



A. Herlert                                  ISCC November 16, 2009

				
DOCUMENT INFO
Shared By:
Categories:
Tags:
Stats:
views:2
posted:7/23/2012
language:
pages:19